OpenStudy (anonymous):

How do you... Evaluate P(-1/2) if P(x) = 7x4 + 7x2 - 4

OpenStudy (anonymous):

you have to plug in -1/2 everywhere you see x in the function

OpenStudy (anonymous):

Like this : P(-1/2)=7(-1/2)^4+7(-1/2)^2-4

OpenStudy (anonymous):

then you have to solve it?

OpenStudy (anonymous):

Then just find out what the numbers are

OpenStudy (anonymous):

and add them

OpenStudy (anonymous):

do you have a calculator

OpenStudy (anonymous):

some kind of TS thing that I don't really understand how to use

OpenStudy (anonymous):

well, type the equation in exactly like i wrote it out and you will get the answer

OpenStudy (anonymous):

you still there?

OpenStudy (anonymous):

yea, still trying to figure it out because my granddaughter seems to have absconded somewhere with my calculator

OpenStudy (anonymous):

well, then right click on what i wrote copy it and plug it into google, you will get the answer

OpenStudy (anonymous):

ok, I think I used the wrong google, I got some reeeeeeeeeealy strange stuff.....lol

OpenStudy (gw2011):

P(-1/2)=7(-1/2)^4+7(-1/2)^2-4 P(-1/2)=7(1/16)+7(1/4)-4 P(-1/2)=7/16+7/4-4 P(-1/2)=7/16+28/16-64/16 P(-1/2)=-29/16

OpenStudy (anonymous):

OMG, tytytytyty so much

OpenStudy (anonymous):

I'm never going to finish this assignment....lol

OpenStudy (gw2011):

You're welcome